# Mergers of compact objects with cores of massive stars: evolution, r-process and multi-messenger signatures

### v1

@@ Abstract
The study of massive binary systems has steadily progressed over the past decades, with increasing focus on their evolution, interactions and mergers, driven by improvements in computational modelling and observational techniques. In particular, when a binary system involves a massive giant and a neutron star (NS) or a black hole (BH) that go through common envelope evolution (CEE), it might results in the merger of the compact object with the core of its giant companion, giving rise to various high energy astrophysical phenomena. We review the different evolutionary channels that lead to compact objects-core mergers, key physical processes with emphasis on the role of accretion physics, feasibility of r-process nucleosynthesis, expected observable electromagnetic, neutrinos and gravitational-waves (GWs) signatures, as well as potential correlation with detected core collapse supernovae (CCSNe), luminous fast blue optical transients (LFBOTs) and low luminosity long gamma-ray bursts (LGRBs). After presenting our current understanding of these mergers, we conclude discussing prospects for future advancements.
@@ 1 INTRODUCTION
Compact objects are frequently found in binary systems. When a neutron star (NS) or a black hole (BH) is close enough to a massive companion, the immense swelling of the companion as it evolves might result in the engulfment of the compact object that starts spiralling inside its envelope due to dynamical friction. This evolutionary phase, which is commonly known as common envelope evolution (CEE; e.g., Paczynski 1976; Taam et al. 1978; Iben & Livio 1993; Izzard et al. 2012; Ivanova et al. 2013; Röpke & De Marco 2023), could result either in the ejection of the envelope and survival of the compact object-core binary (e.g., Vigna-Gómez et al. 2018), or their merger within the envelope gas (e.g., Grichener 2023). The vast accretion by the compact object as it merges with the core of the giant could lead to the formation of an accretion disk (e.g., Papish et al. 2015; Grichener & Soker 2019; Everson et al. 2024). In cases where jets launched by this accretion disk play a key role in the merger, the resultant transient is termed common envelope jets supernova (CEJSN; Soker & Gilkis 2018).
Mergers of NSs/BHs with the cores of their giant companions have been attracting interest as potential r-process nucleosynthesis sites (e.g., Papish et al. 2015; Grichener & Soker 2019; Grichener et al. 2022; Grichener 2023; Jin & Soker 2024) and potential progenitors of long gamma-ray bursts (LGRBs; e.g., Fryer & Woosley 1998; Zhang & Fryer 2001; Thöne et al. 2011; Rueda & Ruffini 2012; Fryer et al. 2014), merger driven core collapse supernovae (CCSNe; e.g., Barkov & Komissarov 2011; Chevalier 2012; Schrøder et al. 2020; Soker & Gilkis 2018; Gofman & Soker 2019; Dong et al. 2021) and luminous fast blue optical transients (LFBOTs; e.g., Soker et al. 2019; Metzger 2022; Soker 2022; Cohen & Soker 2023). Due to their complexity and the wide variety of physical processes they encompass, many studies perform semi-analytical and one-dimensional simulations to explore these mergers (e.g., Soker et al. 2019; Grichener & Soker 2019; Grichener & Soker 2021; Cohen & Soker 2023), while others simulate selected aspects in three dimensions (e.g., Schrøder et al. 2020; Schreier et al. 2021; Hillel et al. 2022a; Hillel et al. 2022b; Everson et al. 2024). However, to date, computational barriers do not allow for fully self-consistent simulations of these transients. The aim of this review is to present our current understanding of NS/BH-core mergers and explore the role they play in high energy astrophysics, while highlighting the limitations of the currently performed studies, and presenting future directions that are essential for addressing the open questions that remain in this field.
@@ 2 Evolutionary channels
The mergers of NSs and BHs with cores of giant stars are believed to have a significant role in explaining luminous transients (section 5.1). Understanding the formation channels of binary systems that result in these mergers is important for determining where they are most likely to occur and identify precursor signatures that might assist in their detection. Grichener (2023) performs binary population synthesis to study the possible evolutionary routes of binary systems towards NS/BH-core mergers and predict their rates. They identify two distinct evolution channels leading to NS-core mergers (Fig. 1), which correspond to different regimes in the initial orbital separation distribution of the progenitor binary system (left panel of Fig. 2; orange bins). Only the main evolution channel (left panel of Fig. 1) can lead to the merger of a BH with the core of a giant star, which is associated with the single mode in the right panel of Fig. 2 (turquoise bins). Fig. 2 shows the initial orbital separation distribution of system that results in NS/BH core mergers for a population synthesis model with a common envelope efficiency parameter (Livio & Soker, 1988) of and solar metallicity, but the trends shown in this plot are robust for all the studied parameters (see Grichener 2023 for more details).
The most dominant channel (left panel of Fig. 1), which was previously discussed by Chevalier (2012) and Schrøder et al. (2020), begins with a binary system that is composed of two massive stars relatively close to one another (initial orbital separation of ). The initially more massive star (hereafter referred to as the primary star) evolves past the main sequence (MS), fills its Roche lobe and begins transferring mass to its lighter MS companion (hereafter referred to as the secondary star). This dynamically stable mass transfer episode might lead to the rejuvenation of the secondary star (e.g., Hellings 1984; Renzo et al. 2023; Wagg et al. 2024), modifying the core-envelope boundary region due to the accreted mass and decreasing significantly the envelope binding energy for the rest of the evolution. The mass transfer remains stable and the primary eventually gets stripped off its envelope due to the mass lost to the secondary accretor (e.g., Kippenhahn & Weigert 1967; Götberg et al. 2017; Götberg et al. 2018) or through stellar winds (e.g., Maeder 1981. See Beasor & Smith 2022 for a different view), leaving a naked core that explodes as a stripped-envelope supernova (SN; e.g., Laplace et al. 2021; Vartanyan et al. 2021), observationally defined as type IIb or Ib/Ic, and produces an NS or a BH remnant.
@@
Alternatively, if the tidal disruption radius lies within the core of the secondary star, the compact object would enter while the core still possesses its original structure. In this case the mass accretion could occur either through an accretion disk formed by vast accretion onto the compact object as it spirals through the core, leading to a bright transient (e.g., Soker & Gilkis 2018; Soker et al. 2019) or be quasi-spherical powering a Thorne-Zytkow object (TZO) in case of an NS primary (e.g., Thorne & Zytkow 1975; Thorne & Zytkow 1977. See O’Grady et al. 2024 for a recent review) or a quasi-star in case of a BH primary (e.g., Begelman et al. 2006; Begelman et al. 2008; Ball et al. 2011), depending on the degree to which the core was spun-up by the compact object during CEE. If the specific angular momentum of the core material, , is larger than the angular momentum needed to maintain an orbit at the innermost stable circular orbit around a compact object, , then an accretion disk will immediately form and a stable TZO configuration will be prevented, as found in the stellar evolution and hydrodynamical simulations analyzed by Everson et al. (2024). Nathaniel et al. (2024) implement the condition of the tidal disruption radius in population synthesis data of massive binaries to explore a wider parameter space and determine which systems will avoid disruption and form classical TZOs.
If the entire envelope is ejected before the compact object reaches the proximity of the core then they will avoid merging, and the stripped core might explode in a second stripped-envelope SN, leaving a second compact object behind. In cases where the compact objects in the double compact object (DCO) binary system remain close enough after the natal-kick induced by the SN explosion, they might merge within Hubble time producing GWs that are detectable with current facilities (e.g., Tutukov & Yungel’Son 1993; Chevalier 2012; Vigna-Gómez et al. 2018; Grichener 2023). Table 1 summarizes the main evolutionary outcomes of a giant star - compact object CEE system.
@@ 3 Accretion and feedback physics during the merger
where is the density near the compact object. For a NS that mergers with an evolved core of a massive star, this accretion rate can reach a value of several hundredths solar masses per second (e.g., Grichener & Soker 2019).
This analytical prescription does not take into account the possible effects of density gradients present in the envelope and core of the giant, tidal disruption of the core by the compact object, and ejection of gas by jets launching on the accretion rate. Therefore, conducting three dimensional hydrodynamical simulations to model this accretion is necessary for obtaining more accurate results. Hydrodynamical simulations that compute the accretion rates by a companion during different stages of CEE disagree on whether it constitutes a considerable fraction of the Bondi-Hoyle-Lyttleton mass accretion (e.g., Livio et al. 1986; Chamandy et al. 2018; López-Cámara et al. 2020a; Kashi et al. 2022; Prust et al. 2024) or if it is much lower (e.g., Ricker & Taam 2012; MacLeod & Ramirez-Ruiz 2015a; MacLeod & Ramirez-Ruiz 2015b). The different results are due to the high complexity in modelling CEE involving compact objects, that does not allow to include all relevant physical processes (such as magnetic fields, gravitation and relativistic effects) and forces different studies to make different assumptions.
@@
for a NS and BH accretors, respectively, where is the speed of light, is the radius of the NS, is the typical opacity, is an efficiency factor and is the mass of the BH. For standard parameters, in both cases this accretion rate is on the order of , much lower than the typical Bondi-Hoyle-Lyttleton accretion rate.
In the case of a compact object-core merger, the optically thick in-flowing gas can trap the photons and carry them inwards, leading to adiabatic compression that increases the temperature of the flow. If it reaches the threshold for pair production, that corresponds to the total rest mass energy of the pair , pairs of electrons and positrons can form and annihilate trough the production of photons that get trapped in the flow
@@
which stream out freely due to their low cross section of interaction with matter. The neutrinos escape with a fraction of the accretion energy on a timescale of mili-seconds (Grichener & Soker, 2019), drastically reducing the pressure near the compact object, hence allowing it to accrete at super-Eddington rates (e.g., Houck & Chevalier 1991; Chevalier 1993). 111For BH-core mergers, it is more likely that the gas in-flowing into the BH will carry a larger part of the accretion energy beyond the horizon than the energy lost by neutrinos (e.g., Popham et al. 1999). Esteban et al. (2023) find that detectable MeV neutrino signals are produced during the CEE of an NS with a star due to neutrino cooling that occurs when the NS accretes mass at these high rates. To fully explore the role of neutrino cooling in compact object-core mergers, it is important to note that it can be efficient at increasing the accretion rate by a compact object only in inner regions of the accretion flow where it overpowers viscous heating (e.g., Metzger et al. 2008). In the outer regions viscous heating will drive wind mass loss, reducing the mass accretion rate reaching smaller radii (e.g., Stone & Pringle 2001; Hawley & Balbus 2002; Yuan & Narayan 2014).
@@
If the accretion rate by the compact object is high enough, the specific angular momentum of the accreted gas could be larger than the specific angular momentum of a Keplerian orbit on the surface of the compact object, preventing its free fall onto the compact object and forming a centrifugally-supported accretion disk from the destroyed core material (e.g., Soker 2004). Since the radius of the compact object () is much smaller than the typical orbital separation of the binary system at the onset of the merger (), an accretion disk is expected to be formed. The friction between different segments of the disk cause inner segments to lose angular momentum and fall into the accretor releasing gravitational energy. The energy that is not radiated away could be re-processes to kinetic energy that would speed the rotation of the disk. Energy and angular momentum transport from the disk outwards could be performed through viscosity, magnetic torques and tidal interactions with the compact object (e.g., Pringle & Rees 1972; Shakura & Sunyaev 1973; Dgani et al. 1994), and lead to the launching of disk winds or narrow jets. Even though jets physics is not fully understood, it is generally accepted that magnetic fields play a crucial role in their launching mechanism. Blandford & Payne (1982) propose, for instance, that magnetic fields extract matter from the accretion disk to large distances, where they collimate that gas outflows into opposite narrow jets. The jets can then collide with the surrounding core matter, depositing their kinetic energy into the gas and remain chocked (for more information on the propagation of jets in an external media see section 5.2). This kinetic energy is then converted to thermal energy and radiation, powering a bright transient.
@@
Jets launching could influence the accretion rate by the compact object as it spirals through the envelope and core of the giant via the negative jet feedback mechanism (see Soker 2016 for a review). In this feedback cycle, jets that the accretion disk launches collide with and deposit their kinetic energy in their surroundings gas, expelling some of it away. Since the jets are powered by accretion of matter from the compact object’s surroundings, by expelling the matter around them and leaving less gas to accrete, their power drops. The less powerful jets will then be able to expel less gas around them, until they encounter a fresh supply of new gas as the compact object keeps spiraling-in. Grichener et al. (2021) and Hillel et al. (2022a) find that this negative feedback cycle can lower the accretion rate by a factor of about 10 compared to the accretion rate originally assumed.

### v2

# Mergers of compact objects with cores of massive stars: evolutionary pathways, r-process nucleosynthesis and multi-messenger signatures

@@ Abstract
The study of massive binary systems has steadily progressed over the past decades, with increasing focus on their evolution, interactions and mergers, driven by improvements in computational modelling and observational techniques. In particular, when a binary system involves a massive giant and a neutron star (NS) or a black hole (BH) that go through common envelope evolution (CEE), it might result in the merger of the compact object with the core of its giant companion, giving rise to various high energy astrophysical phenomena. We review the different evolutionary channels that lead to compact object-core mergers, key physical processes with emphasis on the role of accretion physics, feasibility of r-process nucleosynthesis, expected observable electromagnetic, neutrino and gravitational-wave (GW) signatures, as well as potential correlation with detected core collapse supernovae (CCSNe), luminous fast blue optical transients (LFBOTs) and low luminosity long gamma-ray bursts (LGRBs). After presenting our current understanding of these mergers, we conclude discussing prospects for future advancements.
@@ 1 INTRODUCTION
Compact objects are frequently found in binary systems. When a neutron star (NS) or a black hole (BH) is close enough to a massive companion, the immense swelling of the companion as it evolves might result in the engulfment of the compact object that starts spiralling inside its envelope due to dynamical friction. This evolutionary phase, which is commonly known as common envelope evolution (CEE; e.g., Paczynski 1976; Taam et al. 1978a; Iben & Livio 1993; Izzard et al. 2012; Ivanova et al. 2013; Ivanova et al. 2020; Röpke & De Marco 2023), could result either in the ejection of the envelope and survival of the compact object-core binary (e.g., Vigna-Gómez et al. 2018), or their merger within the envelope gas (e.g., Grichener 2023). The vast accretion by the compact object as it merges with the core of the giant could lead to the formation of an accretion disk (e.g., Papish et al. 2015; Grichener & Soker 2019; Everson et al. 2024). In cases where jets launched by this accretion disk play a key role in the merger, the resultant transient is termed common envelope jets supernova (CEJSN; Soker & Gilkis 2018).
Mergers of NSs/BHs with the cores of their giant companions have been attracting interest as potential r-process nucleosynthesis sites (e.g., Papish et al. 2015; Grichener & Soker 2019; Grichener et al. 2022; Grichener 2023; Jin & Soker 2024) and potential progenitors of long gamma-ray bursts (LGRBs; e.g., Fryer & Woosley 1998; Zhang & Fryer 2001; Thöne et al. 2011; Rueda & Ruffini 2012; Fryer et al. 2014), merger driven core collapse supernovae (CCSNe; e.g., Fryer et al. 1996; Barkov & Komissarov 2011; Chevalier 2012; Schrøder et al. 2020; Soker & Gilkis 2018; Gofman & Soker 2019; Dong et al. 2021) and luminous fast blue optical transients (LFBOTs; e.g., Soker et al. 2019; Metzger 2022; Soker 2022; Cohen & Soker 2023). Due to their complexity and the wide variety of physical processes they encompass, many studies perform semi-analytical and one-dimensional simulations to explore these mergers (e.g., Soker et al. 2019; Grichener & Soker 2019; Grichener & Soker 2021; Cohen & Soker 2023), while others simulate selected aspects in three dimensions (e.g., Schrøder et al. 2020; Schreier et al. 2021; Hillel et al. 2022a; Hillel et al. 2022b; Everson et al. 2024). However, to date, computational barriers do not allow for fully self-consistent simulations of these transients. The aim of this review is to present our current understanding of NS/BH-core mergers and explore the role they play in high energy astrophysics, while highlighting the limitations of the currently performed studies, and presenting future directions that are essential for addressing the open questions that remain in this field.
@@ 2 Evolutionary channels
The mergers of NSs and BHs with cores of giant stars are believed to play a significant role in explaining luminous transients (section 5.1). Understanding the formation channels of binary systems that result in these mergers is important for determining where they are most likely to occur and identify precursor signatures that might assist in their detection. Grichener (2023) performs binary population synthesis to study the possible evolutionary routes of binary systems towards NS/BH-core mergers and predict their rates. They identify two distinct evolution channels leading to NS-core mergers (Fig. 1), which correspond to different regimes in the initial orbital separation distribution of the progenitor binary system (left panel of Fig. 2; orange bins). Only the main evolution channel (left panel of Fig. 1) can lead to the merger of a BH with the core of a giant star, which is associated with the single mode in the right panel of Fig. 2 (turquoise bins). Fig. 2 shows the initial orbital separation distribution of system that results in NS/BH core mergers for a population synthesis model with a common envelope efficiency parameter (Livio & Soker, 1988) of and solar metallicity, but the trends shown in this plot are robust for all the studied parameters (see Grichener 2023 for more details).
The most dominant channel (left panel of Fig. 1), which was previously discussed by Chevalier (2012) and Schrøder et al. (2020), begins with a binary system that is composed of two massive stars relatively close to one another (initial orbital separation of ). The initially more massive star (hereafter referred to as the primary star) evolves past the main sequence (MS), fills its Roche lobe and begins transferring mass to its lighter MS companion (hereafter referred to as the secondary star). This dynamically stable mass transfer episode might lead to the rejuvenation of the secondary star (e.g., Hellings 1984; Renzo et al. 2023; Wagg et al. 2024), modifying the core-envelope boundary region due to the accreted mass and decreasing significantly the envelope binding energy for the rest of the evolution. The mass transfer remains stable and the primary eventually gets stripped off its envelope due to the mass lost to the secondary accretor (e.g., Kippenhahn & Weigert 1967; Götberg et al. 2017; Götberg et al. 2018) or through stellar winds (e.g., Maeder 1981. See Beasor & Smith 2022 for a different view), leaving a naked core that explodes as a stripped-envelope supernova (SN; e.g., Podsiadlowski et al. 1992; Laplace et al. 2021; Vartanyan et al. 2021), observationally defined as type IIb or Ib/Ic, and produces an NS or a BH remnant.
@@
Alternatively, if the tidal disruption radius lies within the core of the secondary star and in-spiral is fast enough, the compact object would enter while the core still possesses its original structure. In this case the mass accretion could occur either through an accretion disk formed by vast accretion onto the compact object as it spirals through the core, leading to a bright transient (e.g., Soker & Gilkis 2018; Soker et al. 2019) or be quasi-spherical powering a Thorne-Zytkow object (TZO) in case of an NS primary (e.g., Thorne & Zytkow 1975; Thorne & Zytkow 1977. See O’Grady et al. 2024 for a recent review) or a quasi-star in case of a BH primary (e.g., Begelman et al. 2006; Begelman et al. 2008; Ball et al. 2011), depending on the degree to which the core was spun-up by the compact object during CEE. If the specific angular momentum of the core material, , is larger than the angular momentum needed to maintain an orbit at the innermost stable circular orbit around a compact object, , then an accretion disk will immediately form and a stable TZO configuration will be prevented, as found in the stellar evolution and hydrodynamical simulations analyzed by Everson et al. (2024). Nathaniel et al. (2024) implement the condition of the tidal disruption radius in population synthesis data of massive binaries to explore a wider parameter space and determine which systems will avoid disruption and form classical TZOs.
If the entire envelope is ejected before the compact object reaches the proximity of the core then they will avoid merging, and the stripped core might explode in a second stripped-envelope SN, leaving a second compact object behind. In cases where the compact objects in the double compact object (DCO) binary system remain close enough after the natal-kick induced by the SN explosion, they might merge within Hubble time producing GWs (e.g., Tutukov & Yungel’Son 1993; Chevalier 2012; Vigna-Gómez et al. 2018; Ablimit & Maeda 2018; Grichener 2023) and electromagnetic transients (e.g., Janiuk et al. 2013; Abbott et al. 2017a) that are detectable with current facilities. Table 1 summarizes the main evolutionary outcomes of a giant star - compact object CEE system.
@@ 3 Accretion and feedback physics during the merger
where is the density near the compact object. For a NS that merges with an evolved core of a massive star, this accretion rate can reach a value of several hundredths solar masses per second (e.g., Grichener & Soker 2019).
This analytical prescription does not take into account the possible effects of density gradients present in the envelope and core of the giant, tidal disruption of the core by the compact object, and ejection of gas by jet launching on the accretion rate. Therefore, conducting three dimensional hydrodynamical simulations to model this accretion is necessary for obtaining more accurate results. Hydrodynamical simulations that compute the accretion rates by a companion during different stages of CEE disagree on whether it constitutes a considerable fraction of the Bondi-Hoyle-Lyttleton mass accretion (e.g., Livio et al. 1986; Chamandy et al. 2018; López-Cámara et al. 2020a; Kashi et al. 2022; Prust et al. 2024) or if it is much lower (e.g., Ricker & Taam 2012; MacLeod & Ramirez-Ruiz 2015a; MacLeod & Ramirez-Ruiz 2015b). The different results are due to the high complexity in modelling CEE involving compact objects, that does not allow to include all relevant physical processes (such as magnetic fields, cooling mechanisms, gravitation and relativistic effects) and forces different studies to make different assumptions.
@@
for a NS and BH accretors, respectively, where is the speed of light, is the radius of the NS, is the typical opacity, is an efficiency factor and is the mass of the BH. For standard parameters, in both cases this accretion rate is on the order of , much lower than the typical Bondi-Hoyle-Lyttleton accretion rate. While radiative cooling alone might not be able to increase the accretion rate by compact objects due to the high opacitties of stellar envelopes, convection could assist in carrying energy outwards to regions where radiative transfer becomes more efficient (e.g., Grichener et al. 2018; Wilson & Nordhaus 2019; Wilson & Nordhaus 2022.) To understand the effect of radiation on the accretion process, it is important to account for radiative transport in the simulations, as adiabatic hydrodynamic simulations of Bondi-Hoyle-Lyttleton mass accretion find that the infalling material creates a hydrostatic halo surrounding the compact object (e.g., Ruffert 1994; Prust et al. 2024).
In the case of a compact object-core merger, the optically thick in-flowing gas is likely to trap the photons and carry them inwards, leading to adiabatic compression that increases the temperature of the flow. If it reaches the threshold for pair production, that corresponds to the total rest mass energy of the pair , pairs of electrons and positrons can form and annihilate trough the production of photons that get trapped in the flow
@@
which stream out freely due to their low cross section of interaction with matter. The neutrinos escape with a fraction of the accretion energy on a timescale of mili-seconds (Grichener & Soker, 2019), drastically reducing the pressure near the compact object, hence allowing it to accrete at super-Eddington rates (e.g., Houck & Chevalier 1991; Chevalier 1993; Fryer et al. 1996). 111For BH-core mergers, it is more likely that the gas in-flowing into the BH will carry a larger part of the accretion energy beyond the horizon than the energy lost by neutrinos (e.g., Popham et al. 1999). Esteban et al. (2023) find that detectable MeV neutrino signals are produced during the CEE of an NS with a star due to neutrino cooling that occurs when the NS accretes mass at these high rates. To fully explore the role of neutrino cooling in compact object-core mergers, it is important to note that it can be efficient at increasing the accretion rate by a compact object only in inner regions of the accretion flow where it overpowers viscous heating (e.g., Metzger et al. 2008). In the outer regions viscous heating will drive wind mass loss, reducing the mass accretion rate reaching smaller radii (e.g., Stone & Pringle 2001; Hawley & Balbus 2002; Yuan & Narayan 2014).
@@
If the accretion rate by the compact object is high enough, the specific angular momentum of the accreted gas could be larger than the specific angular momentum of a Keplerian orbit on the surface of the compact object, preventing its free fall onto the compact object and forming a centrifugally-supported accretion disk from the destroyed core material (e.g., Soker 2004). Since the radius of the compact object () is much smaller than the typical orbital separation of the binary system at the onset of the merger (), an accretion disk is expected to be formed. The friction between different segments of the disk cause inner segments to lose angular momentum and fall into the accretor releasing gravitational energy. The energy that is not radiated away could be re-processes to kinetic energy that would speed the rotation of the disk. Energy and angular momentum transport from the disk outwards could be performed through viscosity, magnetic torques and tidal interactions with the compact object (e.g., Pringle & Rees 1972; Shakura & Sunyaev 1973; Dgani et al. 1994), and lead to the launching of disk winds or narrow jets. Even though jet physics is not fully understood, it is generally accepted that magnetic fields play a crucial role in their launching mechanism. Blandford & Payne (1982) propose, for instance, that magnetic fields extract matter from the accretion disk to large distances, where they collimate that gas outflows into opposite narrow jets. The jets can then collide with the surrounding core matter, depositing their kinetic energy into the gas and remain choked (for more information on the propagation of jets in an external media see section 5.2). This kinetic energy is then converted to thermal energy and radiation, powering a bright transient.
@@
Jet launching could influence the accretion rate by the compact object as it spirals through the envelope and core of the giant via the negative jet feedback mechanism (see Soker 2016 for a review). In this feedback cycle, jets that the accretion disk launches collide with and deposit their kinetic energy in their surroundings gas, expelling some of it away. Since the jets are powered by accretion of matter from the compact object’s surroundings, by expelling the matter around them and leaving less gas to accrete, their power drops. The less powerful jets will then be able to expel less gas around them, until they encounter a fresh supply of new gas as the compact object keeps spiraling-in. Grichener et al. (2021) and Hillel et al. (2022a) find that this negative feedback cycle can lower the accretion rate by a factor of about 10 compared to the accretion rate originally assumed.